The average(arithmetic mean) of 6 numbers is 6. If 3 is subtracted from each of 4 of the numbers, what is the new average?
Answer by Theo(13342)

average of 6 numbers is equal to 6
if 3 is subtracted from each of 4 of the numbers, what is the new average?

since the average of the numbers is the result of dividing the total of the numbers by the count of the numbers, then the total of the numbers is the result of multiplying the average of the numbers by the count of the numbers.

for example:

average is 6 and count is 6, therefore total is 6*6 = 36.
total is 36 and count is 6, therefore average is 36/6 = 6.

the total of the 6 numbers is 36 because 6*6 = 36.
take 3 off each of 4 of the numbers and you will be subtracting a total of 3*4 = 12
36 - 12 = 24
the new total is 24, therefore the new average is 24/6 = 4.